# Resource Estimation for Delayed Choice Quantum Entanglement Based Sneakernet Networks Using Neutral Atom qLDPC Memories

### v1

@@ V Results and Discussion
The required E-bit rate ($r_{e}$), or bandwidth, is a critical parameter in our system design. Contemporary communication standards provide context for our bandwidth requirements. For instance, voice communications necessitate 87 kbps, standard definition video conferencing requires 128 kbps, and high-definition video conferencing demands up to 4 Mbps voi . In the realm of text messaging, an average SMS of 120-160 characters consumes approximately 16,000 bits sms , which we consider for our analysis. Note that this is somewhat contrived, but it gives us an anchor to do our analysis. To determine $r_{e}$, we consider several factors: the frequency of one-time pad recharging, the practical time a user might wait at a quantum ATM while recharging ($t_{wait}$), and the average number of bits needed between refills ($b_{req}$). The number of end-users at a destination ($n_{users}$) is constrained by the number of available time slots in a day, calculated as 86,400 seconds divided by $t_{wait}$. Consequently, we express $r_{e}$ as $b_{req}/t_{wait}$.
We estimate $r_{e}$ for SMS usage as an illustrative example. Assuming an average person sends 85 texts daily avg , each requiring 16,000 bits, we calculate $b_{req}$ as $85\times 16,000=1,360,000$ bits. With a $t_{wait}$ of 10 minutes (600 seconds) every day, $r_{e}$ equals 1,360,000/600 = 2.3 kbps or 2.3 KHz.
The gate error rate $p_{g}$ for neutral atom architectures varies from $0.01$ to $0.05$. We take $p_{g}=0.01$ as a near-term realizable optimistic estimate. Unlike most other architectures, neutral atoms have significantly slower 1-qubit gates than 2-qubit gates. Hence we take the 1-qubit gate time $t_{g}=2\mu s$ Wintersperger et al. (2023). Considering the form factor of Pasqal’s processor pas and their claim of producing a $10^{4}$ qubit unit by 2026, we take a rough estimate for our calculations that any truck or train car in the near future can carry a total of $n_{t}=1,000,000$ physical qubits. We also set the number of qATM destinations $S=5$. In this article, we assume intra-metropolitan distances, with typical transport times ranging from one to three hours. We take the teleportation ancilla HGP patch size $n_{a}=\frac{n_{m}}{25}$ because it maintains the same distance as the qLDPC code and the surface code, and its edge is as wide as the size of the logical operators in the qLDPC code Xu et al. (2024).
@@
we can see that to achieve a lower total logical error rate $R_{tot}$, we need larger qLDPC patches. Larger patches have greater code-distance, hence higher tolerance to errors. To achieve greater than 80% fidelity (i.e. $R_{tot}<0.2$), we need patch sizes of at least about 35,000 physical qubits (1400 logical qubits). Patch sizes over 53,000 physical qubits (2120 logical qubits) can achieve fidelities of over 90% while 80,000-qubit-sized patches (3200 logical qubits) can give greater than 95% fidelity. Higher the fidelity, lower is the number of bits wasted for standard reconciliation and privacy amplification, and with fidelities over 95%, this number becomes negligible Bennett et al. (1992). Pasqal, an enterprise dealing with neutral atom quantum architectures, claims that devices with a capacity of 10,000 physical qubits will be a reality in the next couple of years pas , so it would be reasonable to assume larger patch sizes of the order of high $10^{4}$ physical qubits to be achievable in the next few years.
@@
Figure 5: Number of trucks or train cars required for the network, as a function of overall LFR $R_{tot}$, showing an inverse relationship between the network’s tolerable error rate $R_{tot}$ and the required number of transport vehicles (trucks or train cars) for a fixed transit duration of 90 minutes and a fixed bit-rate (bandwidth) $r_{e}$ of 2300 Hz. As the acceptable $R_{tot}$ increases, the fidelity requirement decreases, and smaller patch sizes become viable. These reduced patch sizes correspond to lower code distances, necessitating fewer error correction rounds. Consequently, the unloading process becomes faster, reducing the number of parallel qLDPC patches needed. This cascade effect ultimately results in a decreased demand for transport vehicles to maintain the desired bandwidth.
@@
where $C_{m}$ is the yearly cost of maintaining a single qLDPC memory patch containing $n_{m}$ physical qubits. The numerator represents the yearly cost of maintaining all the qLDPC memory patches in the network, where the factor 2 arises from the fact that we are dealing with Bell-pairs, and $n_{t}/n_{m}$ is the number of quantum computer units on a truck. The denominator represents the number of bits emenated per year by the network. We consider an estimate of $C_{m}=\$2,000,000$ per year per quantum device unit qub . For surface codes, the quantity $n_{m}$ would represent the number of physical qubits in a single quantum device, which would contain multiple surface code patches. We keep $n_{m}$ same for both qLDPC and surface codes.
@@
High number of trucks and consequently high costs per qubit are both mainly attributable to the fact that neutral atom architectures have their physical qubits sparsely spaced in the lattice, and their gate execution times are relatively slow. If we can implement qLDPC codes on faster, more condensed architectures in the future, such as the silicon quantum dot qubits for example Eriksson et al. (2013), we can potentially reduce the number of trucks and consequently the cost per bit by about 3 orders of magnitude.
@@ VI Conclusion
In our analysis, we adopt certain conventions and assumptions based on recent research in quantum error correction codes. The relationship between the number of physical qubits in Low-Density Parity-Check (qLDPC) memory and Surface Code is a key consideration. We assume that a qLDPC memory with $n_{m}$ physical qubits is connected to a surface code with $n_{m}/25$ physical qubits for teleportation. This ratio is consistent with the simulations presented in ref. Xu et al. (2024), although it is not explicitly stated as a requirement. While it is theoretically possible to use surface codes of various sizes, we adhere to this ratio for consistency with existing literature.
Regarding the teleportation scheme for transferring logical qubits between qLDPC and surface codes, our analysis focuses on Hypergraph Product (HGP) codes. RefXu et al. (2024) specifically defines this teleportation scheme for HGP codes, but not for Lifted Product (LP) codes. While it may be possible to extend this scheme to LP codes, such an extension would require additional assumptions about the ability to generate logical X operators at the edge of the patch. To maintain rigor and avoid unsubstantiated claims, we restrict our analysis to HGP codes, for which the teleportation scheme is well-defined. This decision ensures that our results are based on established protocols and avoids potential inaccuracies that could arise from extrapolating beyond the current state of knowledge in the field. Designing a teleportation protocol for these other non-HGP qLDPC code families can reduce patch sizes by a factor of two to four, reducing our resource requirements.
Another caveat is that of classically transferring measurement results onto the users’ smartphones. In future research, if a protocol can be devised to teleport the logical qubits of a quantum error-correcting code patch onto a single photon, then photon measurement device can be built into smartphones, which makes the transfer quantum-secure. An alternate way, as outlined in the pseudocode 2 is to perform a heralded qubit transfer from each atomic physical qubit on the surface code to a photon Kalb et al. (2015); Covey et al. (2023) and measure those photons on the end-user’s smartphone Lowndes et al. (2021) in a pre-defined sequence and decode to extract the logical surface-code measurement, but this method has only 0.88 fidelity and 0.69 efficiency Kalb et al. (2015), which when propagated over all the atoms of a single surface code, would lead to high failure rates and high resource overheads. Therefore, development of a more efficient method of performing atom-to-photon qubit transfer would be beneficial towards a highly secure network with minimal resource overhead.

### v2

@@ V Results and Discussion
The required E-bit rate ($r_{e}$), or bandwidth, is a critical parameter in our system design. Contemporary communication standards provide context for our bandwidth requirements. For instance, voice communications necessitate 87 kbps, standard definition video conferencing requires 128 kbps, and high-definition video conferencing demands up to 4 Mbps [43]. In the realm of text messaging, an average SMS of 120-160 characters consumes approximately 16,000 bits [44], which we consider for our analysis. Note that this is somewhat contrived, but it gives us an anchor to do our analysis. To determine $r_{e}$, we consider several factors: the frequency of one-time pad recharging, the practical time a user might wait at a quantum ATM while recharging ($t_{wait}$), and the average number of bits needed between refills ($b_{req}$). The number of end-users at a destination ($n_{users}$) is constrained by the number of available time slots in a day, calculated as 86,400 seconds divided by $t_{wait}$. Consequently, we express $r_{e}$ as $b_{req}/t_{wait}$.
We estimate $r_{e}$ for SMS usage as an illustrative example. Assuming an average person sends 85 texts daily [45], each requiring 16,000 bits, we calculate $b_{req}$ as $85\times 16,000=1,360,000$ bits. With a $t_{wait}$ of 10 minutes (600 seconds) every day, $r_{e}$ equals 1,360,000/600 = 2.3 kbps or 2.3 KHz.
The gate error rate $p_{g}$ for neutral atom architectures varies from $0.001$ to $0.005$. Quantinuum, a quantum computing company, has demonstrated a one-qubit gate fidelity of $0.999979$ and a two-qubit gate fidelity of $0.99914$ [46]. Therefore it would be reasonable to consider $p_{g}=1-0.9992=0.0008$ as a near-term realizable optimistic estimate.
Idling Errors: Idling errors can be modeled by adjusting the two-qubit gate error rate $p_{g}$ as follows [27]:
$p_{g}\rightarrow p_{g}\left(1+\frac{3t_{r}(n)}{0.005T_{c}}\right)$ (20)
where $T_{c}=10$s is the coherence time of neutral atom qubits. If we plug in a worst-case scenario of a patch size of $100,000$, we see that $p_{g}$ gets scaled up by a factor of $1.2754$ i.e. by $27.54\%$. Considering our original estimate of $p_{g}=0.0008$, the new $p_{g}$ becomes $p_{g}=0.0008*1.2754=0.001$, which we use for our analysis.
Unlike most other architectures, neutral atoms have significantly slower 1-qubit gates than 2-qubit gates. Hence we take the 1-qubit gate time $t_{g}=2\mu s$ [40]. Considering the form factor of Pasqal’s processor [47] and their claim of producing a $10^{4}$ qubit unit by 2026, we take a rough estimate for our calculations that any truck or train car in the near future can carry a total of $n_{t}=1,000,000$ physical qubits. We also set the number of qATM destinations $S=5$. In this article, we assume intra-metropolitan distances, with typical transport times ranging from one to three hours. We take the teleportation ancilla HGP patch size $n_{a}=\frac{n_{m}}{25}$ because it maintains the same distance as the qLDPC code and the surface code, and its edge is as wide as the size of the logical operators in the qLDPC code [27].
@@
we can see that to achieve a lower total logical error rate $R_{tot}$, we need larger qLDPC patches. Larger patches have greater code-distance, hence higher tolerance to errors. To achieve greater than 80% fidelity (i.e. $R_{tot}<0.2$), we need patch sizes of at least about 35,000 physical qubits (1400 logical qubits). Patch sizes over 53,000 physical qubits (2120 logical qubits) can achieve fidelities of over 90% while 80,000-qubit-sized patches (3200 logical qubits) can give greater than 95% fidelity. Higher the fidelity, lower is the number of bits wasted for standard reconciliation and privacy amplification, and with fidelities over 95%, this number becomes negligible [48]. Pasqal, an enterprise dealing with neutral atom quantum architectures, claims that devices with a capacity of 10,000 physical qubits will be a reality in the next couple of years [47], so it would be reasonable to assume larger patch sizes of the order of high $10^{4}$ physical qubits to be achievable in the next few years.
@@
Figure 5: Number $N_{truck_{tot}}$ of trucks or train cars required for the network, as a function of overall LFR $R_{tot}$, showing an inverse relationship between the network’s tolerable error rate $R_{tot}$ and the required number of transport vehicles (trucks or train cars) for a fixed transit duration of 90 minutes. As the acceptable $R_{tot}$ increases, the fidelity requirement decreases, and smaller patch sizes become viable. These reduced patch sizes correspond to lower code distances, necessitating fewer error correction rounds. Consequently, the unloading process becomes faster, reducing the number of parallel qLDPC patches needed. This cascade effect ultimately results in a decreased demand for transport vehicles to maintain the desired bandwidth.
@@
where $C_{m}$ is the yearly cost of maintaining a single qLDPC memory patch containing $n_{m}$ physical qubits. The numerator represents the yearly cost of maintaining all the qLDPC memory patches in the network, where the factor 2 arises from the fact that we are dealing with Bell-pairs, and $n_{t}/n_{m}$ is the number of quantum computer units on a truck. The denominator represents the number of bits emenated per year by the network. We consider an estimate of $C_{m}=\$2,000,000$ per year per quantum device unit [49]. For surface codes, the quantity $n_{m}$ would represent the number of physical qubits in a single quantum device, which would contain multiple surface code patches. We keep $n_{m}$ same for both qLDPC and surface codes.
@@
Nevertheless, even with qLDPC codes, the cost of \$1.40 per qubit would mean an average of \$22,400 per text message at 16,000 bits per message, which is extremely expensive, thus making the protocol in its current form commercially infeasible. High number of trucks and consequently high costs per qubit are both mainly attributable to the fact that neutral atom architectures have their physical qubits sparsely spaced in the lattice, and their gate execution times are relatively slow. If we can implement qLDPC codes on faster, more condensed architectures in the future, such as the silicon quantum dot qubits for example [50], we can potentially reduce the number of trucks and consequently the cost per bit by about 3 orders of magnitude.
@@ VI Conclusion
In our analysis, we adopt certain conventions and assumptions based on recent research in quantum error correction codes. The relationship between the number of physical qubits in Low-Density Parity-Check (qLDPC) memory and Surface Code is a key consideration. We assume that a qLDPC memory with $n_{m}$ physical qubits is connected to a surface code with $n_{m}/25$ physical qubits for teleportation. This ratio is consistent with the simulations presented in ref. [27], although it is not explicitly stated as a requirement. While it is theoretically possible to use surface codes of various sizes, we adhere to this ratio for consistency with existing literature.
Regarding the teleportation scheme for transferring logical qubits between qLDPC and surface codes, our analysis focuses on Hypergraph Product (HGP) codes. Ref[27] specifically defines this teleportation scheme for HGP codes, but not for Lifted Product (LP) codes. While it may be possible to extend this scheme to LP codes, such an extension would require additional assumptions about the ability to generate logical X operators at the edge of the patch. To maintain rigor and avoid unsubstantiated claims, we restrict our analysis to HGP codes, for which the teleportation scheme is well-defined. This decision ensures that our results are based on established protocols and avoids potential inaccuracies that could arise from extrapolating beyond the current state of knowledge in the field. Designing a teleportation protocol for these other non-HGP qLDPC code families can reduce patch sizes by a factor of two to four, reducing our resource requirements.
Another caveat is that of classically transferring measurement results onto the users’ smartphones. In future research, if a protocol can be devised to teleport the logical qubits of a quantum error-correcting code patch onto a single photon, then photon measurement device can be built into smartphones, which makes the transfer quantum-secure. An alternate way, as outlined in the pseudocode 2 is to perform a heralded qubit transfer from each atomic physical qubit on the surface code to a photon [51, 52] and measure those photons on the end-user’s smartphone [31] in a pre-defined sequence and decode to extract the logical surface-code measurement, but this method has only 0.88 fidelity and 0.69 efficiency [51], which when propagated over all the atoms of a single surface code, would lead to high failure rates and high resource overheads. Therefore, development of a more efficient method of performing atom-to-photon qubit transfer would be beneficial towards a highly secure network with minimal resource overhead.